OpenStudy (anonymous):

strontium-90 has a half-life of 28 years how long will it take for all strontium 90 to presently be reduced by 1/32nd

OpenStudy (joannablackwelder):

There are multiple ways to do this problem, but one pretty straightforward way is to notice that you will have to divide the original amount in half 5 times to be left with 1/32 of what you started with.

OpenStudy (joannablackwelder):

Original amount/2=amount left after 1 half life

OpenStudy (joannablackwelder):

Amount left after 1 half life/2= amount left after 2 half lives

OpenStudy (joannablackwelder):

Notice that if you substitute original amount/2 for amount left after 1 half life, you would get (Original amount/2)/2=amount left after 2 half lives

OpenStudy (joannablackwelder):

Dividing by 2 twice is the same as dividing by 2 squared.

OpenStudy (joannablackwelder):

So, original amount / (2^2) = amount left after 2 half lives

OpenStudy (joannablackwelder):

Similarly, Original amount /(2^5) = amount left after 5 half lives

OpenStudy (joannablackwelder):

Also notice that 2^5=32